OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve for (f ⋅ g)(2) and select the correct answer below: f(x) = 3x2 + 2 g(x) = x − 8 −42 −84 42 84

OpenStudy (anonymous):

@iGreen

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi :D

OpenStudy (misty1212):

\[f(x)=3x^2+2\] right?

OpenStudy (anonymous):

Yeah

OpenStudy (misty1212):

ok then ' \[f(2)=3\times 2^2+2\] what is that number?

OpenStudy (anonymous):

I got 68

OpenStudy (misty1212):

think you did the "order 'o operations" wrong square first, then multiply

OpenStudy (anonymous):

How do I square?

OpenStudy (igreen):

\(\sf 2^2 \rightarrow 2 \times 2\)

OpenStudy (misty1212):

how do you square a number? multiply it by itself so \[2^2=2\times 2\]

OpenStudy (anonymous):

so it'd be f(2)=3×2x2+2?

OpenStudy (misty1212):

yes what is \[ 3\times 2\times 2+2\]?

OpenStudy (anonymous):

14

OpenStudy (misty1212):

right now hold on to that number

OpenStudy (anonymous):

Alright

OpenStudy (misty1212):

\[g(x)=x-8\] what is \[g(2)\]?

OpenStudy (anonymous):

g(2) = x - 8?

OpenStudy (misty1212):

hmmm i see you are a bit confused as to how to evaluate a function

OpenStudy (anonymous):

Very :/

OpenStudy (misty1212):

\[g(x)=x-8\\ g(\spadesuit)=\spadesuit-8\\ g(\xi)=\xi -8\]

OpenStudy (misty1212):

so how do you find \(g(2)\)? where you see an \(x\) replace it by a \(2\)

OpenStudy (anonymous):

g(2) = 2 - 8? The other two functions have clovers and & signs in them, not sure if it's meant to be there.

OpenStudy (misty1212):

yeah\[g(2)=2-8\]

OpenStudy (anonymous):

So what's the next step?

OpenStudy (misty1212):

find \[2-8\]

OpenStudy (anonymous):

-6, if i'm solving from left to right

OpenStudy (anonymous):

g(2) = -6?

OpenStudy (misty1212):

yes it is \(-6\) no matter what you do

OpenStudy (misty1212):

final step \[f\times g(2)=f(2)\times g(2)=14\times (-6)\]

OpenStudy (anonymous):

so the final answer woulds be -84?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

Thanks :)) Could you help me with a few more?

OpenStudy (misty1212):

lol sure we can at least do one more

OpenStudy (anonymous):

Sweet :)

OpenStudy (anonymous):

Gaming systems are on sale for 20% off the original price (g), which can be expressed with the function p(g) = 0.8g. Local taxes are an additional 12% of the discounted price (p), which can be expressed with the function c(p) = 1.12p. Using this information, which of the following represents the final price of a gaming system with the discount and taxes applied? c(p) + p(g) = 1.92g c[p(g)] = 0.896g g[c(p)] = 1.92p c(p) ⋅ p(g) = 0.896pg

OpenStudy (anonymous):

This seems like an easy one, it's just the trouble of setting up the function I guess

OpenStudy (misty1212):

it is easy it is the last one

OpenStudy (anonymous):

How did you get that?

OpenStudy (misty1212):

lets do it with an item that cost $100

OpenStudy (misty1212):

first take the 20% discount, multiply \[100\times .8=80\]

OpenStudy (misty1212):

so it is $80 before the tax then put in the tax get \[80\times 1.12=89.6\]

OpenStudy (misty1212):

what did we do? first multiply by \(.8\) then multiply by \(1.12\)

OpenStudy (misty1212):

ohh hold the phone for a second

OpenStudy (misty1212):

i see what they want you to say, they want you to pick B

OpenStudy (anonymous):

Oooh so it's B?

OpenStudy (misty1212):

first do \(p\) then do \(c\) it is \[c(p(g))\]

OpenStudy (misty1212):

which is kind of silly since you can do it just by multiplying, but whatever pick B

OpenStudy (anonymous):

Alright that's what I was thinking too haha, thank you. Couple more :D?

OpenStudy (anonymous):

Or would you rather me open a new question to give you another best response?